OpenStudy (anonymous):

C = 17(0.9x) + 13(0.05x) + 0(0.05x) What is the minimum number of forms she must complete in a month to earn a total of at least $1,276?

OpenStudy (anonymous):

can you please specify what 'x' and C represent? I assume 'x' is the number of forms and C is her earning, but just being sure.

OpenStudy (anonymous):

Based on her regular performance, the equation below can be used to determine C, the total amount Rachel earns a month entering data, where x represents the total number of forms she completes that month.

OpenStudy (anonymous):

Basically you just need to expand your initial equation and then solve for x so you will get C = 15.3x + 0.65x + 0 C = 15.95x You know C, so you have 15.95x = 1276 Divide both sides by 15.95 and you get x = 1276/15.95 x = 80 She needs to fill 80 forms

OpenStudy (anonymous):

oh i didnt unstand it thanx